how do you solve this?Find x
2x+4=8

Found 2 solutions by Alan3354, Jessica789:
Answer by Alan3354(69443)

Find x
2x+4=8
=========
Subtract 4
Divide by 2.
---------------
Or,
Divide by 2 first, then .... WDYT?

Answer by Jessica789(1)

First subtact 4 from both sides of the equation which gives you
2x+4-4=8-4
Then from doing that you get
2x=4
Then divide both sides of the equation by 2 which give you
2x/2=4/2
Which simplifies to
X=2